Find the ๐‘›th order differential coefficient of ๐‘ฆ = sin ๐‘ฅ log(๐‘Ž๐‘ฅ + ๐‘).

1
Expert's answer
2022-03-16T14:25:19-0400

"y=sinx log(ax+b)"

Using product rule, we have

"\frac{dy}{dx}(UV)=U\frac{dv}{dx}+V\frac{du}{dx}"

"\frac{dy}{dx}=sinx\frac{d}{dx}log(ax+b)+log(ax+b)\frac{d}{dx}sinx"


"=sinx(\frac{1\times a}{ax+b})+log(ax+b)cosx"

"\frac{d^2y}{dx}=" "sinx\frac{d}{dx}(\frac{1\times a}{ax+b})+(\frac{1\times a}{ax+b})\frac{d}{dx}sinx+
log(ax+b)cosx+cos\frac{d}{dx}log(ax+b)"


"=sinx\frac{(-1)^1\times a^2}{(ax+b)^2}+\frac{1\times a}{ax+b}cosx+log(ax+b)(-sinx)

+cosx\frac{(-1)^1\times a^2}{(ax+b)^2}"

"\frac{d^3y}{dx}=sinx\frac{(-1)^2\times a^3\times2}{(ax+b)^3}
+\frac{(-1)^1\times a^2}{(ax+b)^2}cosx+\frac{1\times a}{ax+b}(-sinx)
+cosx\frac{(-1)^1\times a^2}{(ax+b)^2}+log(ax+b)(-cosx)
+cosx\frac{(-1)^2\times a^3\times 2}{(ax+b)^3}"


"=sinx\frac{-1^2\times a^3\times 2}{(ax+b)^2}+2(cosx\frac{-1^1\times a^2}{(ax+b)^2}
+log(ax+b)(-cosx)+cosx\frac{(-1)^2\times a^3\times 2}{(ax+b)^3}"

"\frac{d^4y}{dx}=sinx\frac{-1^3\times a^4\times 6}{(ax+b)^4}

+\frac{-1^2\times a^3\times 2}{(ax+b)^3}cosx
+\frac{-1^1\times a^2}{(ax+b)^2}(-sinx)+\frac{-1^2\times a^3\times 2}{(ax+b)^3}cosx
+\frac{1\times a}{ax+b}(-cosx)+(-sinx)\frac{-1^1\times a^2}{(ax+b)^4}
+cosx\frac{-1^2\times a^3\times 2}{(ax+b)^3}
+\frac{-1^1\times a^2}{(ax+b)^2}(-sinx)+
log(ax+b)sinx+(-cosx)\frac{1\times a}{ax+b}+cosx\frac{-1^3\times a^4\times 6}{(ax+b)^4}
+(-sinx)\frac{-1^2\times a^3\times 2}{(ax+b)^3}"

"=sinx\frac{-1^3\times a^4\times 6}{(ax+b)^4}+3(cosx\frac{-1^2\times a^3\times 2}{(ax+b)^3})
-3(sinx\frac{-1^1\times a^2}{(ax+b)^2})-2(cosx\frac{1\times a}{ax+b})
+log(ax+b)sinx
-sinx\frac{-1^2\times a^3\times 2}{(ax+b)^3}+cosx\frac{-1^3\times a^4\times 6}{(ax+b)^4}"


"\frac{d^ny}{dx}=" "sinx\frac{-1^{n-1}(n-1){!}a^n}{(ax+b)^{n}})+(n-1)(cos x\frac{-1^{n-2}(n-2){!}a^{n-1}}{(ax+b)^{n-1}})
-(n-1)(sinx\frac{-1^{n-3}(n-3){!}a^{n-2}}{(ax+b)^{n-2}})
-(n-2)(cos x\frac{1\times a}{ax+b})
+log(ax+b)(sin(x+\frac{n\pi}{2})-(sinx\frac{-1^{n-2}(n-2){!}a^{n-1}}{(ax+b)^{n-1}})+cos x\frac{-1^{n-1}(n-1){!}a^n}{(ax+b)^{n}})"